OpenStudy (anonymous):

If a power x =b power y= ab, then find the value of y. i) (x-1)/x ii) (x+1)/x iii) x/(x-1)

OpenStudy (anonymous):

pls rewrite the first line with commas to separate things n make them clear

OpenStudy (anonymous):

a^x=b^y=ab, then find "y" in terms of "x" as given in the options.

OpenStudy (anonymous):

I have done like this. Plz say whether it is correct or not. Given, a^x=ab Or, a^(x-1)=b Also, given that, b^y=ab Substituting, we get, [a^(x-1)]^y=ab Or, a^(xy-y)=a^x Or, xy-y=x Or, y(x-1)=x Or, y = x/(x-1)

OpenStudy (anonymous):

let me check....

OpenStudy (anonymous):

its good.... u r right....

OpenStudy (anonymous):

Thanks a lot..